# Supermassive black hole growth from stellar binary encounters

### v2

@@ 5 Conclusion
This analysis explores SMBH growth from accretion of stars captured through binary break-up. We presented a simple theoretical framework to assess the effectiveness of this growth channel (§2). We then applied it to 30 galaxies hosting SMBHs with well-constrained masses, predicting both past and future growth by this mechanism (§3). In most cases, this growth channel can account for at least 10% of the black hole mass acquired over the past Gigayear. In 20% of the galaxies, 50% or more of their mass was obtained in this way.
Our model, an extension of one proposed by Bromley et al. (2012), has its limitations. These include uncertainties in the demographics and intrinsic properties of binary stars in the central regions of galaxies, including whether the orbital distribution of binaries is consistent with a full loss cone (Lightman & Shapiro, 1977), as we have assumed here. A comparison with data from HVSs that originate from interaction with the LMC’s black hole (§4) is a promising indicator that some of these uncertainties may be reduced in the future. Broader comparisons with tidal disruption events and their rate of occurrence (e.g., Stone & Metzger, 2016) may also inform our model.
@@ Acknowledgments
We thank C. Hannah for generously providing the SMBH data that formed the foundation of this analysis, as well as guidance from A. Seth. AJ is grateful for support from the University of Utah’s Undergraduate Research Opportunity Program.

### v3

@@ 6 Conclusion
We present a simple theoretical framework to assess SMBH growth from accretion of stars captured through binary break-up (§2). Our approach is based on kinematical rates, including a simple $nv\sigma$” model and a modified version designed to accommodate loss-cone effects and limited reservoirs of stellar mass. We apply our models to a set of nearby galaxies hosting SMBHs, predicting both past and future growth by this mechanism (§3). A comparison between tidal disruption event rates and our approach (§4) suggests that our modified kinematical model realistically tracks the encounter rate of binary stars, providing some confidence in our predictions. Finally, in §5, we examine black hole growth in the LMC, following the recent analysis of Han et al. (2025) and observed hypervelocity stars that directly reveal binary break-ups.
Our results argue that binary break-up and captured star accretion can play a significant role in black hole growth. Of the 91 galaxies from Hannah et al. (2024), the majority experience a growth rate greater than 10% of the SMBH mass per gigayear in one or the other of our models. Eleven galaxies harbor SMBHs that grow as quickly in both of our models. Our analysis of the putative black hole in the Large Magellanic Cloud, LMC, together with direct evidence for binary break-up (Han et al., 2025), indicates that roughly one third of its mass grew through this channel. Upcoming observations of HVSs in the LMC, together with the anticipated trove of TDE detections with the Vera C. Rubin Observatory (Bricman & Gomboc, 2020), will place meaningful constraints on our predictions.
Our models of black hole growth have limitations. For example, we do not consider the dependence of the fraction of stellar binaries on the stellar mass density or the full phase-space distribution of stars that can interact with an SMBH. We also do not track the bound stars from binary break-up as they accrete onto the black hole. Although we adopt a conservative value for the fraction of binary stars and include a model that compares well with sophisticated numerical calculations of TDE rates (Hannah et al., 2025), missing details may affect inferred SMBH growth rates.
@@ Acknowledgments
We thank C. Hannah for generously providing the SMBH data that formed the foundation of this analysis, as well as guidance from A. Seth. An anonymous referee gave feedback that significantly improved this work. ALJ is grateful for support from the University of Utah’s Undergraduate Research Opportunity Program.